# Date of publication xxxx 00, 0000, date of current version xxxx 00, 0000. xx

### v2

@@ IV-A Step 1: Target Identification
UWB offers high time resolution due to its large bandwidth and narrow pulses. This enables the precise differentiation in time of multiple reflected signals. An illustration of this effect is provided in Fig. 3, which shows the CIRs measured for metal, concrete, and plywood obstacles, overlaid on a background CIR measured without any obstacles. The reflections caused by each material appear as distinct peaks at specific delays relative to the first path.
@@
The QM331x0W chips, which comply with the IEEE 802.15.4z standard, integrate two antennas on the same chip. This standard introduces the STS field mainly for security purposes, but it also enables the creation of additional CIRs that can be used for angle of arrival (AoA) estimation as the same signal reaches each antenna with slightly different phase. On the first antenna, the receiver captures both the preamble CIR and the STS1-based CIR. Then, halfway the UWB packet, the chip switches to the second antenna to capture the STS2-based CIR [single-anchor-localization].
@@
FPFP False positive, the peak that indicates a real object, but is actually noise -
FNFN False negative, the peak that indicates a real object, but is filtered out -
@@ IV-A1 Signal processing
To reliably extract the peak indicating an object, every raw CIR undergoes the following processing phases. Firstly, the CIR samples are initially recorded as I/Q components. To calculate the magnitude of the received signals,  (2) is used on the Preamble CIR and  (3) is used to calculate the phase of the received signals on both STS-based CIRs.
@@
The first four samples of the CIRs are excluded from further processing, as these samples precede the arrival of the first signal and represent the noise floor. However, these initial samples are used later for calculating the Signal-to-Noise-Ratio (SNR) for filtering purposes. At last, to ensure a consistent threshold and comparability across different measurements, the preamble CIR is normalized using min-max scaling.
@@ IV-B Step 2: Filtering
Figure 4: Definition of prominence, width, and the SNR-based score. Prominence is the vertical distance between a peak and its lowest contour line. In this example, the prominence of the potential object’s peak is measured relative to contour line CL1 rather than CL2, since CL2 has a lower value. The width is calculate at half the prominence, using interpolation to determine the intersection points. The SNR-score evaluates the detected peak relative to the noise floor.
@@
  • PDoA: The antenna array’s sensitivity and resolution degrade at wider angles, leading to increased multipath interference and reduced angular accuracy. The calculation of the PDoA, shown in  (5), and adapted from [pdoa-aoa], no longer requires synchronization since both antennas are on the same chip.
@@
The exact values of the filtering parameters and how they impact accuracy will be discussed in Section VI-A.
@@ IV-C2 Angle of Arrival
To estimate the AoA, the PDoA between the two antennas on the same chip is used. This is calculated as shown in  (5), followed by  (7), which shows the calculation of the AoA [pdoa-aoa].
@@ IV-C3 Clustering
By applying the filtering methods from Section IV-B and localising the resulting points relative to the robot using the distance and orientation calculations, a point cloud of potential obstacles is generated. In this final step, a clustering algorithm then removes outliers and groups these points into likely obstacle regions. To this end, the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) is used. DBSCAN is well suited for this work because detected components from real objects tend to appear repeatedly in the same location, forming dense groups of points. By identifying these dense regions, DBSCAN can cluster detected components that originate from the same physical object while simultaneously discarding isolated points as noise. The most important parameter for DBSCAN in this work is epsilon (eps), which defines the maximum distance between two samples for one to be considered in the neighbourhood of the other. The second key parameter is min_samples which is the number of samples in a neighbourhood for a point to be considered as a core point, including the point itself (see Fig. 6). Clusters smaller than these thresholds are discarded as noise. In this paper, eps was set to 20 cm and min_samples to 20 detected peaks. Choosing a smaller value would reduce the number of points labelled as noise, but at the risk of forming incorrect clusters that do not correspond to real objects, which would in turn degrade mapping accuracy. The clustering is performed based on peaks mapped to the Cartesian coordinate system and depends on the number of detected peaks across multiple measurements (these could correspond to different robot positions, although the robot may also remain in the same position (see Fig. 6)).
@@ V Data Collection
The following data collection experiments were conducted to evaluate the effectiveness of the proposed approach and assess its performance under realistic conditions. The data collection experiments were conducted in the controlled Industrial Internet of Things (IIoT) lab at Ghent University [lab-ugent] focusing on object detection and signal behaviour using UWB radar. All experiments were performed using a TurtleBot 4 robot, equipped with one omnidirectional UWB transmitter on the top and multiple dual-antenna, directional UWB receivers on every side depending on the experiment (Fig. 7a). To reduce the impact caused by the first path signal as the transmitter is placed close to the receiver, an absorber is used (Fig. 7b). The specifications of the radar system are listed in Table III. During data capture, the system alternated between channel 5 and channel 9 for each measurement to assess signal behaviour across different frequency bands. The IIoT lab is equipped with a Motion Capture (MOCAP) system, providing millimetre-level accuracy ground truth data as the robot moves.
@@
TABLE III: UWB Radar Specifications - IEEE 802.15.4 compliant
@@
Figure 9: a) First data collection: the robot moves straight forward to the object. The same movement is done three times, once for every material (metal, concrete, plywood). b) Second data collection: two objects are placed, in the front and parallel to the robot while data is captured using two receivers.
@@ VI Results
$\text{{F1-score}}=2\times\frac{\text{P}\times\text{R}}{\text{P}+\text{R}}$ (12)
@@ VI-A Filtering parameters
The three key parameters (peak width, prominence, and SNR-score) were optimised using the dataset obtained from the first data collection (Fig. 9a). A grid search over different parameter values was performed, and each configuration was evaluated using the F1-score and the root mean squared error (RMSE) between the estimated and ground-truth distances. The parameters achieving the highest F1-score and lowest RMSE were selected. To further investigate their behaviour, the relationship between distance and both peak prominence and width was analysed for the three different materials (metal, concrete, plywood). Figures 10a and 10b illustrate the correlations in peak prominence and width as a function of target distance.
Figure 10: Correlation between peak properties and distance: a) Prominence decreases with distance for all materials. b) Width increases as the distance increases for all materials.
As seen in Fig. 10a, peak prominence decreases as the distance increases. At short ranges, the reflected signals retain more energy and produce stronger peaks. With increasing distance, signal attenuation and energy loss cause the peaks to stand out less compared to other peaks or the noise floor. Metal consistently produces the highest peak prominence, which is expected given its high reflectivity. Concrete and plywood, by comparison, show lower prominence because they absorb more energy. At distances lower than 100 cm, concrete cannot be detected with the current setup as the peak prominence and width are too low. As distance grows, both concrete and plywood follow a similar downward trend. This limitation is not exclusive to this material. It can affect any material type depending on its reflectivity. Highly reflective surfaces like metal return more energy producing more prominent peaks even when close to the robot. In contrast, materials such as concrete and plywood absorb more energy, resulting in weaker reflections that are more easily overshadowed by the strong first path signal. When the object is within approximately 60 cm, the reflected peak becomes difficult to distinguish due to the strong first path signal and oversaturation of the receiver [comparative-lidar-radar]. Although an absorber was used between the antennas to mitigate this issue, the results show that the problem persists indicating that antenna design also plays a crucial role. This could be resolved by the design of a dedicated UWB antenna array with better isolation between the individual antenna components [uwb-antenna].
@@
Finally, width, prominence, and SNR-score were tuned using the F1-score, which balances precision and recall. The optimised values are summarised in Table IV, both for specific materials and overall optimal values across different materials. The remainder of this section uses the overall filtering parameters listed in Table IV. These parameters represent the best configuration for mixed-material environments containing metal, concrete, and plywood. If the material in the environment is known beforehand (i.e., only one material type is present), parameters specific to that material type can be used to further improve detection performance.
@@ VI-B1 Impact of different materials after filtering
Figure 11: After step 2: Distance accuracy (errors in cm) for different materials (after filtering on width, prominence, snr-score, and PDoA). Metal is the most reliably detected material in both channel 5 and 9, with an error accuracy lower than 15 cm in 95% of the cases. Concrete and plywood show improved results on channel 9, as the algorithm correctly identifies the correct peak most of the time, with an error accuracy lower than 19 cm in 90% of the cases.
The performance of the algorithm after target identification and filtering (on width, prominence, snr-score, and PDoA) differed across materials and channels, the distance accuracy results are shown in Fig. 11 and the detection probability in Fig. 12. In this step, AoA was not considered, and all detections were assumed in front of the antenna.
Metal: Both channels detected the metallic plate with very high accuracy. On channel 5, the MAE was 6.49 cm with precision and recall above $98\%$. Channel 9 further improved accuracy, reducing the MAE to 5.80 cm and achieving nearly perfect precision and recall.
@@
Overall, channel 9 consistently outperformed channel 5 for non-metallic materials, resulting in lower MAE, fewer outliers, and higher precision/recall. Fig. 11 shows the higher error of concrete and plywood in channel 5 compared to channel 9. In addition, further accuracy gains can be obtained during step 3 of the processing pipeline (see next sections).
(a) Channel 5
(b) Channel 9
Figure 12: After step 2: Detection performance metrics (precision, recall, F1-score) for each material and channel configuration. Metrics are based on correctly retained object peaks (true positives), wrongly kept noise peaks (false positives), and missed object peaks (false negatives) within the range of 20 cm from the ground truth.
VI-B2 Impact of combined filtering and clustering steps
After target identification and filtering using the different methods, the AoA is calculated. The detected peaks are then mapped relative to the robot, followed by clustering. This step further improves accuracy because it removes additional unwanted peaks that correspond to noise, thereby increasing the detection and distance accuracy.
The evolution of the complete algorithm is illustrated in Fig. 13. For this visualisation, data was collected using the parallel antenna in an environment with two objects (illustrated in Fig. 9b). The mobile robot moves in a straight line from start (indicated with a semi-transparent blue circle) to end (indicated with a fully opaque blue circle). Each sub-figure includes a colour bar on the right-hand side representing the SNR-score where a brighter colour indicates higher SNR-score and thus better signal quality. Table V summarises the key parameters used in this work.
TABLE V: Key parameters used in the proposed approach
@@ VI-B2 Impact of combined filtering and clustering steps
width, prominence 1, 0.05 Peak properties: minimum required width and prominence of a peak to be considered significant.
SNR-score 20 Peaks with a SNR-score lower than the threshold are discarded.
AoA_range [45°,45°] Only peaks within this range are kept.
eps 20 cm DBSCAN parameter: max distance between two samples to be part of the same neighbourhood.
min_samples 20 DBSCAN parameter: number of samples in a neighbourhood for a point to be considered as a core point.
min_peaks 50 Minimum number of new peaks before clustering is applied.
range 20 cm Detected peaks within this range are classified as true positives.
bias 15 cm Calibration offset to correct systematic errors in distance measurements.
@@
In the first step (Fig. 13a), a radargram-style visualisation illustrates the robot’s path and the signal quality of the detected peaks, represented by their SNR-score. The data are plotted in an x/y plane to facilitate understanding of subsequent processing steps. In this step no angle of arrival is used, and every detected peak is shown in front of the antenna. At this point, many unwanted peaks caused by noise and unwanted multipath effects are visible, which makes raw CIR unsuitable for direct mapping since it could easily lead to false object detections. To remove the unwanted peaks, the data is filtered in the next stage (Fig. 13b) using width, prominence, and SNR-score defined in Table V. This step significantly reduces unwanted peaks and reveals a clear pattern. It also highlights an important behaviour of the parallel antenna: although it is mounted on the right side of the robot, it still detects peaks created from objects that are not directly in front of it. In this example, peaks from an object located in front of the robot but outside the parallel antenna’s forward-facing range are still recorded, resulting in a line of phantom peaks as shown in the figure diagonally down to the right.

### v3

@@ IV-A Step 1: Target Identification
UWB’s large bandwidth and narrow pulses give it high time resolution, allowing precise separation of multiple reflected signals. Fig. 3 illustrates this by showing the CIRs measured for metal, concrete, and plywood obstacles, overlaid on a background CIR without obstacles. Each material produces distinct reflection peaks at specific delays relative to the first path.
@@
The QM331x0W chips, which comply with the IEEE 802.15.4z standard, integrate two antennas on the same chip. This standard introduces the STS field mainly for security purposes, but it also enables the creation of additional CIRs that can be used for angle of arrival (AoA) estimation as the same signal reaches each antenna with slightly different phase. On the first antenna, the receiver captures both the preamble CIR and the STS1-based CIR. Then, halfway through the UWB packet, the chip switches to the second antenna to capture the STS2-based CIR [21].
@@
FPFP False positive, a peak retained as an object, but actually noise -
FNFN False negative, a peak corresponding to a real object that is filtered out -
@@ IV-A1 Signal processing
To reliably extract the peak indicating an object, every raw CIR undergoes the following processing phases. Firstly, the CIR samples are initially recorded as I/Q components. To calculate the magnitude of the received samples,  (2) is used on the Preamble CIR and  (3) is used to calculate the phase of the received signals on both STS-based CIRs.
@@
The first four samples of the CIRs are excluded from further processing, as these samples precede the arrival of the first signal and represent the noise floor. However, these initial samples are used later for calculating the SNR-based score for filtering purposes. Finally, to ensure a consistent threshold and comparability across different measurements, the preamble CIR is normalized using min-max scaling.
@@ IV-B Step 2: Filtering
Figure 4: Definition of prominence, width, and the SNR-based score. Prominence is the vertical distance between a peak and its lowest contour line. In this example, the prominence of the potential object’s peak is measured relative to contour line CL1 rather than CL2, since CL2 has a lower value. The width is calculated at half the prominence, using interpolation to determine the intersection points. The SNR-score evaluates the detected peak relative to the noise floor.
@@
  • PDoA: The antenna array’s sensitivity and resolution degrade at wider angles, leading to increased multipath interference and reduced angular accuracy. The calculation of the PDoA, shown in  (5), and adapted from [22], no longer requires synchronization since both antennas are on the same chip.
@@ IV-C2 Angle of Arrival
To estimate the AoA, the PDoA between the two antennas on the same chip is used. This is calculated as shown in  (5), followed by  (7), which shows the calculation of the AoA [22].
@@ IV-C3 Clustering
By applying the filtering methods from Section IV-B and localising the resulting points relative to the robot using the distance and orientation calculations, a point cloud of potential obstacles is generated. In this final step, a clustering algorithm then removes outliers and groups these points into likely obstacle regions. To this end, the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) is used. DBSCAN is well suited for this work because detected components from real objects tend to appear repeatedly in the same location, forming dense groups of points. By identifying these dense regions, DBSCAN can cluster detected components that originate from the same physical object while simultaneously discarding isolated points as noise.
The key DBSCAN parameters used are (see Fig. 6):
  • epsilon (eps): the maximum distance between two samples for them to be considered neighbours.
  • min_samples: the number of samples required in a neighbourhood for a point to be considered a core point, including the point itself.
Clusters that do not meet these thresholds are discarded as noise. In this paper, eps was set to 20 cm and min_samples to 20 detected peaks. Using smaller values would reduce the number of points labelled as noise but increases the risk of forming incorrect clusters, which can degrade mapping accuracy. Clustering is performed based on peaks mapped to the Cartesian coordinate system and depends on the number of detected peaks across multiple measurements (these could correspond to different robot positions, although the robot may also remain in the same position (see Fig. 6)).
@@ V Data Collection
The following data collection experiments were conducted to evaluate the effectiveness of the proposed approach and assess its performance under realistic conditions. The data collection experiments were conducted in the controlled Industrial Internet of Things (IIoT) lab at Ghent University [23]. All experiments were performed using a TurtleBot 4 robot, equipped with one omnidirectional UWB transmitter on the top and multiple dual-antenna, directional UWB receivers on every side depending on the experiment (Fig. 7a). Because all antennas are connected to the same onboard computer of the robot, each receiver is synchronised with millisecond accuracy. To reduce the impact caused by the first path signal as the transmitter is placed close to the receiver, an absorber is used (Fig. 7b). The specifications of the radar system are listed in Table III. During data capture, the system alternated between channel 5 and channel 9 for each measurement to assess signal behaviour across different frequency bands. The IIoT lab is equipped with a Motion Capture (MOCAP) system, providing millimetre-level accuracy ground truth data as the robot moves.
@@
TABLE III: UWB Radar Configurations - IEEE 802.15.4 compliant
@@
(a) First data collection
(b) Second data collection
Figure 9: a) The robot moves straight forward to the object. The same movement is done three times, once for every material (metal, concrete, plywood). b) Two objects are placed, in the front and parallel to the robot while data is captured using two receivers.
@@ VI Results
$\text{{F1-score}}=2\frac{\textit{PR}}{P+R}$ (12)
@@ VI-A Filtering parameters
The three key parameters (peak width, prominence, and SNR-score) were optimised using the dataset obtained from the first data collection (Fig. 9a). A grid search over different parameter values was performed, and each configuration was evaluated using the F1-score and the root mean square error (RMSE) between the estimated and ground-truth distances. The parameters achieving the highest F1-score and lowest RMSE were selected. The optimised values are summarised in Table IV, both for specific materials and overall optimal values across different materials. These parameters do not need to be tuned to these exact values. Our analysis shows that a zone exists in the parameter space in which the detection quality remains consistently high, even when the parameters are varied slightly. This stable high-performance zone is noticeably larger for highly reflective materials, whereas less reflective materials such as concrete require parameters closer to their optimal values due to increased sensitivity.
To further investigate their behaviour, the relationship between distance and both peak prominence and width was analysed for the three different materials (metal, concrete, plywood). Figures 10a and 10b illustrate the correlations in peak prominence and width as a function of target distance.
(a) Prominence decreases with distance for all materials
(b) Width increases with distance for all materials
Figure 10: Correlation between peak properties and distance.
As seen in Fig. 10a, peak prominence decreases as the distance increases. At short ranges, the reflected signals retain more energy and produce stronger peaks. With increasing distance, signal attenuation and energy loss cause the peaks to stand out less compared to other peaks or the noise floor. Metal consistently produces the highest peak prominence, which is expected given its high reflectivity. Concrete and plywood, by comparison, show lower prominence because they absorb more energy. At distances lower than 100 cm, concrete cannot be detected with the current setup as the peak prominence and width are too low. As distance grows, both concrete and plywood follow a similar downward trend. This limitation is not exclusive to this material. It can affect any material type depending on its reflectivity. Highly reflective surfaces like metal return more energy producing more prominent peaks even when close to the robot. In contrast, materials such as concrete and plywood absorb more energy, resulting in weaker reflections that are more easily overshadowed by the strong first path signal. When the object is within approximately 60 cm, the reflected peak becomes difficult to distinguish due to the strong first path signal and oversaturation of the receiver [24]. Although an absorber was used between the antennas to mitigate this issue, the results show that the problem persists indicating that antenna design also plays a crucial role. This could be resolved by the design of a dedicated UWB antenna array with better isolation between the individual antenna components [25].
@@
The remainder of this section uses the overall filtering parameters listed in Table IV. These parameters represent the best configuration for mixed-material environments containing metal, concrete, and plywood. If the material in the environment is known beforehand (i.e., only one material type is present), parameters specific to that material type can be used to further improve detection performance.
@@ VI-B1 Impact of different materials after filtering
The performance of the algorithm after target identification and filtering (on width, prominence, SNR-score, and PDoA) differed across materials and channels. The used parameters are illustrated in Table IV. The distance accuracy and detection probability results are shown in Table. V. In this step, AoA was not considered, and all detections were assumed in front of the antenna.
Metal: The metallic plate was detected with consistently high accuracy on both channels. The strong reflectivity produced clear, distinct peaks, resulting in high precision and recall. Channel 9 further improved detection and distance accuracy.
@@
Overall, channel 9 consistently outperformed channel 5 for non-metallic materials, resulting in lower MAE, fewer outliers, and higher precision/recall. In addition, further accuracy gains can be obtained during step 3 of the processing pipeline (see next sections).
TABLE V: After step 2: Distance accuracy (in cm) and detection probability (%) after filtering. MAE denotes the mean absolute error, SD the standard deviation, and P90/P95 the 90th and 95th percentiles of the absolute distance error, respectively.
Distance Accuracy (in cm) Detection Probability (%)
Channel Material MAE P90 P95 SD Precision Recall F1-score
5 Metal 6.49 12.34 12.34 14.25 98.93 100.00 99.46
Concrete 19.92 44.24 116.62 36.96 81.25 93.77 87.06
Plywood 30.74 48.02 231.51 63.08 79.30 68.04 73.24
9 Metal 5.80 14.33 14.34 4.70 99.67 100.00 99.83
Concrete 10.50 18.49 33.32 11.37 90.71 88.40 89.54
Plywood 7.99 18.81 18.82 7.40 98.14 99.23 98.68
TABLE VI: After step 3: Distance accuracy (in cm) and detection probability (%) after filtering, angle of arrival, and mapping. MAE denotes the mean absolute error, SD the standard deviation, and P90/P95 the 90th and 95th percentiles of the absolute distance error, respectively.
Distance Accuracy (in cm) Detection Probability (%)
Channel Material MAE P90 P95 SD Precision Recall F1-score
5 Metal 8.21 16.63 23.76 7.49 93.20 99.93 96.44
Concrete 9.01 18.43 24.18 6.82 93.73 91.55 92.63
Plywood 11.24 19.71 29.65 8.63 93.31 79.26 85.71
9 Metal 6.85 14.73 17.40 5.87 96.93 99.11 98.01
Concrete 16.89 39.19 46.47 13.90 73.42 83.38 78.08
Plywood 8.47 18.87 18.90 6.96 98.55 99.10 98.82
VI-B2 Impact of combined filtering and mapping
TABLE VII: Key parameters used in step 3 (filtering + mapping).
@@ VI-B2 Impact of combined filtering and mapping
AoA_range [45°,45°] Valid AoA interval.
eps 20 cm Max distance between two samples to be part of the same neighbourhood (DBSCAN).
min_samples 20 Number of samples in a neighbourhood for a point to be considered as a core point (DBSCAN).
min_peaks 50 Minimum number of new peaks before clustering.
range 20 cm True positive distance range.
bias ch5 - 15 cm Distance calibration offset.
bias ch9 - 13 cm
bias_AoA ch5 - 0.0522 rad Angular calibration offset.
bias_AoA ch9 - 0.0209 rad
After AoA estimation, the detected peaks are no longer assumed to be located directly in front of the robot. Instead, their positions are calculated relative to the robot and mapped onto the x/y plane. The UWB chips used in the proposed approach have an angular measurement accuracy of approximately 5°. As a result, differences can occur between the true and estimated angle of arrival, depending on the channel and material used, as explained later in this section. At larger distances, even small AoA estimation errors result in significant position errors.
The filtering parameters from Table IV and mapping parameters from Table VII are used. Table VI shows that the detection performance remains largely unchanged after this step. Precision and recall slightly increase for plywood, while a decrease is observed for concrete on channel 9, with an improvement on channel 5. This confirms that AoA accuracy is influenced by both channel selection and material properties. A clear reduction in standard deviation (SD) is observed across nearly all channel-material combinations, indicating a lower spread in the error distribution. This confirms that the clustering step effectively removes spatially inconsistent detections that do not belong to a valid cluster. In general, the remaining distance errors stay below 30 cm, except for concrete on channel 9, where larger AoA deviations lead to increased errors.
VI-B3 Complete pipeline visualisation
The evolution of the complete algorithm is illustrated in Fig. 11. For this visualisation, data were collected using the parallel antenna on channel 5 in an environment with two objects (illustrated in Fig. 9b). The filtering parameters from Table IV and mapping parameters from Table VII are used. The mobile robot moves in a straight line from start (indicated with a semi-transparent blue circle) to end (indicated with a fully opaque blue circle). Each sub-figure includes a colour bar on the right-hand side representing the SNR-score where a brighter colour indicates higher SNR-score and thus better signal quality.
@@ VI-B3 Complete pipeline visualisation
In the first step (Fig. 11a), a radargram-style visualisation illustrates the robot’s path and the signal quality of the detected peaks, represented by their SNR-score. The data are plotted in an x/y plane to facilitate understanding of subsequent processing steps. In this step no angle of arrival is used, and every detected peak is shown in front of the antenna. At this point, many unwanted peaks caused by noise and unwanted multipath effects are visible, which makes raw CIR unsuitable for direct mapping since it could easily lead to false object detections. To remove the unwanted peaks, the data is filtered in the next stage (Fig. 11b) using width, prominence, and SNR-score. This step significantly reduces unwanted peaks and reveals a clear pattern. It also highlights an important behaviour of the parallel antenna: although it is mounted on the right side of the robot, it still detects peaks created from objects that are not directly in front of it. In this example, peaks from an object located in front of the robot but outside the parallel antenna’s forward-facing range are still recorded, resulting in a line of phantom peaks as shown in the figure diagonally down to the right.